# The role of porosity in the transition to inertial regime in porous media flows

### v2

@@
1 Introduction
Many real-life flows through porous media occur at velocities at which the effects caused by fluid inertia cannot be omitted. Such scenarios are of much importance for gas extraction [1, 2], mixing and reactive processes [3, 4], or flow through network-like porous media [5], e.g. the circulatory system [6, 7]. It is thus crucial from the scientific and application point of view to well well understand the physical principles of the transition of flow through porous media from the creeping flow (Darcy) regime to the inertial regime and the behavior of the flow within the latter. Of special interest are the relations between the pore-scale hydrodynamics and the effective (volume-averaged) hydrodynamic coefficients of a porous medium, which serve as a link connecting the fundamental fluid dynamical processes to the more practical quantities characterizing those systems.
@@ 1 Introduction
$\boldsymbol{U}+\beta\rho\boldsymbol{U}|\boldsymbol{U}|=\frac{k_{0}}{\nu}\left(-\langle\nabla P\rangle+\rho\boldsymbol{g}\right)$ (2)
where $|\boldsymbol{U}|$ is the magnitude of the superficial velocity, $\beta$ is the free coefficient and the porous medium is again assumed to be isotropic. The transition from the Darcy to the inertial regime is often associated with the significant reorganization of the velocity and pressure fields within the pore spaces. Among many aspects of such reorganization, it was observed that the fluid kinetic energy becomes more evenly distributed in the pore space [13], which has to do with significant mass flow through an increasing number of channels and confinement of more kinetic energy within the vortices [14, 15, 16]. Our recent studies on high- [14] and low-porosity [16] porous media suggest that the porosity is one of the factors determining the dynamics of the transition from Darcy to inertial regime. Nevertheless, the exact mechanism of how the geometry of the pores impacts the onset of inertia is still unknown.
@@
2 Methods
2.1 Physical and numerical setup
2.1.1 Governing equations and boundary conditions
@@
The acceleration is aligned with the $x$-axis, $\boldsymbol{g}=[g,0,0]$, and $\boldsymbol{n}$ is the local unit normal vector on the solid boundary. We adopt the boldcase notation for the vectors and refer to the Cartesian coordinates with indices $0,1,2$, e.g. $\boldsymbol{V}\equiv[V_{0},V_{1},V_{2}]$. We denote the magnitude of a vector with a non-boldcase font, e.g. $|\boldsymbol{V}|\equiv V$.
@@
2.1.2 Stochastic porous medium
@@
To solve the governing equations in the stochastic samples, we use Lattice Boltzmann Method (LBM) [30, 31] implemented for the GPU execution using Template Numerical Library [32, 33]. Due to the uniform discretization of space, LBM allows for simple treatment of complex boundaries and has been successfully used to model flows in porous media [34, 35, 36], We use D3Q19 velocity set for $\phi=0.7$ sample and D3Q27 set for $\phi=0.9$ sample, for better stability at higher velocities. Two-relaxation time (TRT) collision kernel [37] with $\Lambda=1/4$ and the symmetric relaxation time $\tau^{+}=0.505$ is used, along with the lattice speed of sound $c_{s}=1/\sqrt{3}$. The discretization in space consists of $N=304$ lattice sites along each direction for $\phi=0.7$ sample and $N=160$ lattice sites along each direction for $\phi=0.9$ sample. The no-slip walls are implemented via the halfway-bounceback method [30, 31]. In the simulations with the highest Reynolds number, in the final timestep, the maximal value of either velocity component for either sample did not exceed $3.73\cdot 10^{-2}$ and the density variation $|1-\rho|$ did not exceed $1.73\cdot 10^{-3}$, all in lattice units.
The steady-state is assumed to be reached when the relative difference between the mass fluxes along $x_{0}$-direction per unit time measured at $x_{0}=0.25$ defined as
@@
and measured every $10^{4}$ timesteps falls below $10^{-2}$. In the above equation, $Q_{0}^{t}$ denotes the mass flux along $x_{0}$-direction at time $t$.
@@
Figure 2: Visualization of the geometries of the solid phase of the two stochastic porous samples considered in this work. The porosities are $\phi=0.7$ (left) and $\phi=0.9$ (right). The black-edged cube denotes the boundaries of the periodic cell.
2.1.3 Simple cubic porous medium
The two simple cubic (SC) porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We consider porosities $\phi=0.1$ and $\phi=0.999$ corresponding to the obstacle radii $r=0.6526$ and $r=0.062$, respectively. For the convenience of visualization and meshing, the computational geometry for the lower-porosity sample considers the spherical obstacles placed in each of the periodic cell’s corners, and for the higher-porosity sample, we place a single obstacle in the center of the domain (see Fig. 3).
For SC samples, we solve the governing equations with openFOAM [38]. We use the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) algorithm [39] for the pressure-velocity coupling, implemented in simpleFOAM solver. In contrast to LBM used with the stochastic samples, SIMPLE allows one to solve directly for the steady-state solution, shortening the calculation time. For $\phi=0.1$, the convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. All considered systems reach this convergence criterion, and the verification of the steady-state solution for the system of the highest flow rate is performed with pimpleFOAM. The space discretization is generated by first creating the regular cubic grid with the discretization parameter $h=1/195$ using blockMesh and then adjusting it to the actual fluid volume geometry with snappyHexMesh (see Fig. 4).
For $\phi=0.999$, the convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. Three systems of the highest flow rate did not reach this criterion; however, the transient simulations performed with pimpleFOAM allow them to reach the steady state, and those data are used in the subsequent analysis. The determination of the steady state in those cases is based on the relative difference between the current and the estimated final value of the $x_{0}$-component of velocity in the bulk and in the wake of the obstacle. In the final iteration, this difference does not exceed $10^{-3}$. The final values of the velocities were estimated from the fit of the exponential function to the values of those velocities in time. We note that the $\phi=0.999$ system becomes unsteady in the range of $Re_{k}\in\>(5\cdot 10^{3};10^{4}]$. The discretization for $\phi=0.999$ systems solved using simpleFOAM is created the same way as for the $\phi=0.1$ systems. For the three highest flow rates solved with pimpleFOAM, we first create a regular cubic grid with the discretization parameter $h_{\text{bulk}}=1/75$ using blockMesh and then adjust it to the actual fluid volume geometry with snappyHexMesh with local refinement ($h_{\text{min}}=1/150$) near the sphere and in its wake (see Fig. 4). For all transient simulations (for both porosities), we choose the timestep $\Delta t$ such that the maximal Courant number in the system is below unity.
@@ 2.1.3 Simple cubic porous medium
Figure 3: Visualization of the geometries of the solid (gray) and fluid (blue, opaque) phases of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (left) and $\phi=0.999$ (right). The black-edged cube denotes the boundaries of the periodic cell. In the case of the low-porosity sample, only a few chosen obstacles residing in the periodic cell are shown for clarity.
@@
Figure 4: Visualization of the meshes of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (two top subplots) and $\phi=0.999$ (two bottom subplots, for the uniform and refined discretization). The top-right subplot shows the magnified view of the top-left mesh. The black-edged cube denotes the boundaries of the periodic cell. Only $z<0.5$ half of each mesh is shown.
2.2 The used definitions of the tortuosity
@@
Figure 5: Schematic representation of a streamline $s$ as a curve tangent to the fluid velocity field vectors (black arrows) in each of the streamline’s points (two chosen are shown here as black dots, $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$). An infinitesimal increment along a streamline, $d\boldsymbol{s}$ (small blue arrow), is parallel to the local velocity vector, here normalized for convenience. Any of the streamline’s points identifies it uniquely ($s(\boldsymbol{x})\equiv s(\boldsymbol{x^{\prime}})$).
@@
with $\lambda(\boldsymbol{x})$ denoting the (curvilinear) length of the streamline, $ds$ the length of the infinitesimal increment along the streamline, and $L_{\text{ref}}$ denoting the reference length. Unless otherwise stated, we take the reference length to be equal to the length of the porous sample’s side, i.e., $L_{\text{ref}}=L$. The integration in Eq. (5) is performed over the intersection of a plane $A$ perpendicular to the mean flow direction and the fluid volume $\Omega$ (see Fig. 6, leftmost plot). Also, $A$ must be intersected by all the streamlines in the system. As it is ambiguous to tell the length of a looping streamline inside a recirculation zone, the integration in Eq. (5) is usually limited only to the intersection of the plane perpendicular to the mean flow direction and the percolating volume of the fluid (see Fig. 6, middle plot)
@@
where the integration goes over the whole fluid volume $\Omega$ (see Fig. 6, rightmost plot). We will refer to this definition as the volume-integrated tortuosity. The notation $I_{\alpha}[\>\cdot\>]$ is a shorthand for the integral of some quantity over the fluid volume or its subset, i.e.
@@
Figure 6: Schematic representation of the three definitions of tortuosities from Eqs. (5), (7), and (8). $T_{s}^{*}$ (left) and $T_{s}$ (middle) integrate over the surface elements $d\boldsymbol{x}$ of the intersections of an arbitrary surface AA and the corresponding fluid volumes. $T_{s}^{*}$ considers all the fluid volume $\Omega$, while $T_{s}$ only its percolating part, $\Omega_{p}$. $T_{\Omega}$ (right) integrates over the whole fluid volume $\Omega$. $T_{s}$ and $T_{\Omega}$ are used further in the present study.
2.3 Results analysis
2.3.1 Generation of streamlines
To generate the streamlines in SC samples, we numerically integrate the advection equation for the massless tracers advected in the fluid velocity field. We use the first-order explicit Euler method for the timestepping and meshless RBF-FD approximation [40] for the interpolation in space. Cubic radial functions augmented with second-order monomials are chosen as the interpolation basis, along with 25 closest neighbors in the approximation stencil. We use the C++ implementation of meshless approximation routines provided by Medusa library [41]. For the generation of streamlines, we use the normalized velocity field, $\boldsymbol{V}/V$, i.e., we set the magnitude of the velocity vectors to unity and preserve only their direction (see Fig. 7).
Figure 7: Schematic representation of the streamline generation procedure. The current increment of the streamline, $\Delta\boldsymbol{s}$ (denoted with an arrow), starts with the interpolation of the velocity to the current streamline end (blue filled circle). The interpolation uses a meshless scheme (and is thus naturally capable of operating on non-regular grids) and considers a number of the mesh points closest to the current streamline tip (filled squares, encircled in a dashed loop). The next streamline point (open circle) is calculated using the first-order Euler time integration.
The initial positions of the tracers depend on the task at hand. Due to the very different geometries and steady-state velocity fields in each of the SC samples, we tailor-cut the streamline generation for each case. For $\phi=0.1$ SC sample, for the calculation of the streamline-based tortuosity $T_{s}$, we place them uniformly at the fluid part of the $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers along $x_{1}$- and $x_{2}$-direction (see Fig. 8). To find the recirculation/percolating volumes, we place the tracers only in the vicinity of the solid boundaries at $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers. For this task, the grid from which the positions of the tracers are taken is rotated by $\pi/4$ about $x_{0}$-axis for a better compliance with the solid boundaries. For both tasks, we use the timestep length $\Delta t_{s}=10^{-3}$ during the integration along characteristics. Considering the initial positions of the streamlines for both tasks, theoretically, none of them should hit the solid wall, have a negative increment in $x$-direction, and all the streamlines should connect the inlet and the outlet (should be percolating). If any streamline fails to satisfy the mentioned conditions, we assume this is due to numerical errors and exclude it from further analysis.
Figure 8: Initial positions of the streamlines for the calculation of $T_{s}$ (left) and the identification of the recirculation/percolating volume (middle) for the low-porosity SC sample. The rightmost image shows which part of the domain is shown in the remaining ones. The fluid volume $\Omega$ is depicted in blue.
In case of $\phi=0.999$ SC sample, for the calculation of $T_{s}$ we place the tracers on a square grid of step $h_{s}=10^{-2}$ at the $x_{0}=0$ plane (see Fig. 9, top-left) and use the timestep length $\Delta t_{s}=10^{-2}$. We dismiss the streamlines that do not fulfill the previously stated criteria. Due to numerical errors in the solution of the governing equations, the mean values of $x_{1}$- and $x_{2}$-velocities in this case are not identically zero, although they are two orders of magnitude lower than the mean $x_{0}$-velocity. It does not impact the calculation of $T_{\Omega}$ significantly, however, it causes the nearly-straight streamlines to experience a drift in a direction perpendicular to $x_{0}$. This results in their elongation form $L$ by a certain value, the magnitude of which is comparable to the value of $T_{s}-1$. In turn, the streamline-based tortuosity is artificially increased to a non-negligible extent. Due to this, for the higher-porosity SC sample, in Eq. (6) we use the distance between the streamline’s start and end as the normalization length, rather than the porous sample’s side length $L$. For the determination of the recirculation/percolating volume, we use two sets of the initial positions of the tracers (see Fig. 9, top-right). To properly capture the flow separation on the rear part of the sphere’s surface, we place the tracers upwind from the sphere, in the vicinity of the domain’s centerline $x_{1}=x_{2}=0.5$ (see Fig. 9, bottom-left). After the integration, we apply the same criteria to dismiss certain streamlines as previously. To properly capture the saddle point at the end of the recirculation zone in the wake of the sphere, we place the tracers downwind from the sphere (also in the vicinity of the centerline) and integrate them in the upwind direction (see Fig. 9, bottom-right). We do not apply the previously described streamline exclusion criteria; rather,, we stop the integration before the streamlines reach the separation ring on the sphere’s surface, to avoid them falling into the recirculation zone due to the numerical errors. We use the separation $h_{s}=5\cdot 10^{-4}$ between the streamline starting positions along $x_{1}$- and $x_{2}$-direction and the timestep length $\Delta t_{s}=10^{-3}$ for the identification of the recirculation/percolating volume.
Figure 9: Initial positions of the streamlines for the calculations in the high-porosity SC sample. Top-left: for the calculation of $T_{s}$. Top-right: for the identification of recirculation/percolating volumes. Bottom row: magnified views at the initial positions of the forward- (A) and backward-integrated (B) streamlines. The $x_{1}=x_{2}=0.5$ line is shown for clarity.
2.3.2 Determining the recirculation and percolating volumes of fluid
To determine which points of the fluid domain $\Omega$ of the SC samples belong to the recirculation or percolating volume, the crucial part is to approximate the boundary between the two, $\partial\Omega_{pv}$. To do this, first we divide the domain into $N_{H}^{3}$ cubic cells and consider only those occupied by the fluid. We use $N_{H}=160$ for both systems. We then mark those of the cells which contain at least one point of the streamlines generated for the purpose of identification of $\Omega_{p}$ and $\Omega_{v}$ (Fig. 10, left). As mentioned above, we generate only the percolating streamlines (although not all of them). Due to this, holes may appear in the clusters of marked cells, denoted by an asterisk. Having obtained the cluster of the marked cells, we consider those lying on the cluster’s boundary (i.e. having at least one non-marked neighbor), near the expected position of $\partial\Omega_{pv}$, as the approximation of $\partial\Omega_{pv}$ (Fig. 10, middle). Lastly, all the cells lying on either side of the approximate boundary are marked as belonging to the recirculation or the percolating volume. The last step of this procedure automatically takes care of the proper coloring of the cells which do not contain any streamlines but should do so, if sufficiently many percolating streamlines were generated (the ones marked with an asterisk in Fig. 10). This way, one needs to generate the streamlines only in the vicinity of $\partial\Omega_{pv}$.
Figure 10: Schematic representation of the procedure for identifying the percolating ($\Omega_{p}$) and recirculation ($\Omega_{v}$) volumes. Each square represents a cell into which we divide the fluid volume. We note that those cells do not comply with the discretization. The percolating streamlines are plotted with thick dashed lines. We refer the reader to the paragraph Determining the recirculation and percolating volumes of fluid in Sec. 2.3 for the detailed explanation of the identification procedure.
2.3.3 Calculation of $T_{\Omega}$ constituents
3 Results
3.1 Difference in tortuosity–Reynolds number relation for porous samples of different porosities.
We first investigate the transition from the Darcy to inertial regime for two stochastic porous samples of porosities $\phi=0.7$ and $\phi=0.9$. Left subplot of Fig. 11 shows the dependence of the tortuosity $T_{\Omega}$ defined in Eq. (8) on the Reynolds number
$Re_{k}=\frac{\sqrt{k(Re_{k}\rightarrow 0)}\langle V_{0}\rangle\phi}{\nu},\quad\langle V_{0}\rangle=\frac{I[V_{0}]}{I[1]}$ (11)
where $k(Re_{k}\rightarrow 0)$ is the permeability at the vanishing Reynolds number. The same data are shown rescaled in the right subplot of Fig. 11, where by $T_{\Omega}(Re_{k}\rightarrow 0)$ we mean the value of $T_{\Omega}$ at the vanishing Reynolds number (denoted by the horizontal dashed line in the left subplot). We note that for $Re_{k}$ values beyond the investigated range, the solution went unsteady. We observe that in the Stokes regime, the tortuosity for each sample has a constant value, $1.07$ for $\phi=0.9$ and $1.175$ for $\phi=0.7$. As $Re_{k}$ increases, the tortuosity first decreases and then increases for both samples. However, the magnitude of the initial decrease and subsequent increase is considerably different for the two porosities. In the case of the higher-porosity sample, the difference between the minimum of $T_{\Omega}$ and the $T_{\Omega}(Re_{k}\rightarrow 0)$ is about $0.02$ ($0.3$ relative to the Darcy-regime value above unity, see the right subplot), and the increase is about $0.01$. For the lower-porosity sample, these values are $0.003$ ($0.01$ relative to the Darcy-regime value above unity) and $0.125$ ($0.67$ relative to the Darcy-regime value above unity), respectively. Such differences between the minimal and maximal values achieved by $T_{\Omega}$ in the steady-state regime were previously observed in stochastic high-porosity media [14] and in low-porosity geological samples [16]. The onset of inertial effects is also visible in the visualizations of the velocity fields, as the weaker channeling at the outlets of the samples (Fig. 12) and the emergence of the recirculation zones in the wakes of the obstacles (Fig. 13).
Figure 11: Left: the values of $T_{\Omega}$ for stochastic porous media of porosity $\phi=0.7$ (circles) and $\phi=0.9$ (squares) during the Darcy-inertial transition. The horizontal dashed lines denote the values of $T_{\Omega}$ in the limit of vanishingly small $Re_{k}$ ($Re_{k}\rightarrow 0$). Right: the same data as in the left subplot but rescaled by $T_{\Omega}$ at vanishingly small $Re_{k}$.
3.2 Splitting the volume-integrated tortuosity $T_{\Omega}$
To explain such a different behavior of $T_{\Omega}(Re_{k})$ in the low- and high-porosity stochastic systems, we first take a closer look at the definition of $T_{\Omega}$ and observe that both the numerator and denominator can be split into separate contributions coming from the percolating volume $\Omega_{p}$ and the recirculation volume $\Omega_{v}$
@@ 3.2 Splitting the volume-integrated tortuosity $T_{\Omega}$
According to the definition of the vortex from Sec. 2, the mean velocity of the vortex in steady-state flow is zero, i.e., $I[\boldsymbol{V}]_{v}=\boldsymbol{0}$, so the term $I[V_{0}]_{v}$ in the denominator can be omitted. Thus, we can rewrite Eq. (12) as
@@
Such a form of the volume-integrated tortuosity allows one to clearly distinguish between the contributions from the percolating and the recirculation zones and directly points to the sources of inequality in Eq. (10). The denominator, $I[V_{0}]_{p}$, is proportional to the Reynolds number (cf. Eq. (11)) and does not introduce any new unknown variable into our considerations.

### v3

@@
I Introduction
Many real-life flows through porous media occur at velocities at which the effects caused by fluid inertia cannot be omitted. Such scenarios are of much importance for gas extraction [1, 2], mixing and reactive processes [3, 4], or flow through network-like porous media [5], e.g. the circulatory system [6, 7]. It is thus crucial from the scientific and application point of view to well understand the physical principles of the transition of flow through porous media from the creeping flow (Darcy) regime to the inertial regime and the behavior of the flow within the latter. Of special interest are the relations between the pore-scale hydrodynamics and the effective (volume-averaged) hydrodynamic coefficients of a porous medium, which serve as a link connecting the fundamental fluid dynamical processes to the more practical quantities characterizing those systems.
@@ I Introduction
$\boldsymbol{U}+\beta\rho\boldsymbol{U}|\boldsymbol{U}|+{\color[rgb]{0,0,0}\gamma\boldsymbol{U}|\boldsymbol{U}|^{2}}=\frac{k_{0}}{\nu}\left(-\langle\nabla P\rangle+\rho\boldsymbol{g}\right)$ (2)
where $|\boldsymbol{U}|$ is the magnitude of the superficial velocity, $\beta$ and $\gamma$ are the free coefficients and the porous medium is again assumed to be isotropic. The transition from the Darcy to the inertial regime is often associated with the significant reorganization of the velocity and pressure fields within the pore spaces. Among many aspects of such reorganization, it was observed that the fluid kinetic energy becomes more evenly distributed in the pore space [13], which has to do with significant mass flow through an increasing number of channels and confinement of more kinetic energy within the vortices [14, 15, 16]. Our recent studies on high- [14] and low-porosity [16] porous media suggest that the porosity is one of the factors determining the dynamics of the transition from Darcy to inertial regime. Nevertheless, the exact mechanism of how the geometry of the pores impacts the onset of inertia is still unknown.
@@
II Methods
II.1 Physical and numerical setup
II.1.1 Governing equations and boundary conditions
@@
The acceleration is aligned with the $x$-axis, $\boldsymbol{g}=[g,0,0]$, and $\boldsymbol{n}$ is the local unit normal vector on the solid boundary. We adopt the boldcase notation for the vectors and refer to the Cartesian coordinates with indices $0,1,2$, e.g. $\boldsymbol{V}\equiv[V_{0},V_{1},V_{2}]$. We denote the magnitude of a vector with a non-boldcase font, e.g. $|\boldsymbol{V}|\equiv V$. In the actual simulations, we set $\rho=1$ and $\nu=1$. We note that all the quantities in Eqs. (3) are given in non-dimensionalized form and such will be presented in the further part of this work. To provide a better perspective on the real-life values of the solution fields in the studied systems, we provide exemplary conversion factors to the dimensionalized forms of the pressure and velocity in Appendix C.1.
@@
II.1.2 Stochastic porous medium
@@
To solve the governing equations in the stochastic samples, we use Lattice Boltzmann Method (LBM) [30, 31] implemented for the GPU execution using Template Numerical Library [32, 33]. Due to the uniform discretization of space, LBM allows for simple treatment of complex boundaries and has been successfully used to model flows in porous media [34, 35, 36], We use D3Q19 velocity set for $\phi=0.7$ sample and D3Q27 set for $\phi=0.9$ sample, for better stability at higher velocities. Two-relaxation time (TRT) collision kernel [37] with $\Lambda=1/4$ and the symmetric relaxation time $\tau^{+}=0.505$ is used, along with the lattice speed of sound $c_{s}=1/\sqrt{3}$. The discretization in space consists of $N=304$ lattice sites along each direction for $\phi=0.7$ sample and $N=160$ lattice sites along each direction for $\phi=0.9$ sample. The no-slip walls are implemented via the halfway-bounceback method [30, 31]. In the simulations with the highest Reynolds number, in the final timestep, the maximal value of any velocity component for either sample did not exceed $3.73\cdot 10^{-2}$ and the density variation $|1-\rho|$ did not exceed $1.73\cdot 10^{-3}$, all in lattice units. We provide a more detailed description of the used LBM model in Appendix C.2.
The steady-state is assumed to be reached when the relative difference between the mass fluxes along $x_{0}$-direction per unit time measured every $10^{4}$ timesteps at $x_{0}=0.25$ defined as
@@
falls below $10^{-2}$. In the above equation, $Q_{0}^{t}$ denotes the mass flux along $x_{0}$-direction at time $t$.
@@
Figure 2: Visualization of the geometries of the solid phase of the two stochastic porous samples considered in this work. The target porosities are $\phi=0.7$ (left) and $\phi=0.9$ (right). The black-edged cube denotes the boundaries of the periodic cell.
II.1.3 Simple cubic porous medium
The four simple cubic (SC) porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We consider porosities $\phi=0.1,\>0.59,\>0.93,\>0.999$ corresponding to the obstacle radii r=0.6526, 0.46, 0.26, 0.062r=0.6526,\>0.46,\>0.26,\>0.062, respectively. For the convenience of visualization and meshing, the computational geometry for the lowest-porosity sample considers the spherical obstacles placed in each of the periodic cell’s corners, and for the other samples, we place a single obstacle in the center of the domain (see Fig. 3). We solve the governing equations using SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) algorithm [38] implemented in simpleFOAM solver of openFOAM [39] and use regular discretization with the mesh parameter $h=1/195$ for $\phi=0.1$ (see Fig. 4, top row) and $h=1/150$ for the remaining samples (see Fig. 4, bottom-left for the highest-porosity sample). In contrast to LBM, SIMPLE allows to directly solve for the steady-state, reducing the simulation time. We use pimpleFOAM solver to confirm the steady state in several simulations of the highest Reynolds number. In the case of the highest-porosity sample, we do so on variable-density meshes with the target cell size ranging from $h_{\text{bulk}}=1/75$ to $h_{\text{min}}=1/150$ (see Fig. 4, bottom-right). We include a more detailed description of the numerical setup for SC samples in Appendix C.C.3. We provide the analysis of the solution convergence with the refined mesh parameter ($h$ or $h_{\text{bulk}}$ and $h_{\text{min}}$) for SC systems in Appendix B. For all transient simulations, we choose the timestep $\Delta t$ such that the maximal Courant number in the system is below unity.
@@ II.1.3 Simple cubic porous medium
Figure 3: Visualization of the geometries of the solid (gray) and fluid (blue, opaque) phases of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1,\>0.59,\>0.93,\>0.999$, counting from the top-left to the bottom-right. The black-edged cube denotes the boundaries of the periodic cell. In the case of the lowest-porosity sample, only a few chosen obstacles residing in the periodic cell are shown for clarity.
@@
Figure 4: Visualization of the meshes of two chosen simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (two top subplots) and $\phi=0.999$ (two bottom subplots, for the uniform and refined discretization). The top-right subplot shows the magnified view of the top-left mesh. The black-edged cube denotes the boundaries of the periodic cell. Only $z<0.5$ half of each mesh is shown.
II.2 The used definitions of the tortuosity
@@
Figure 5: Schematic representation of a streamline $s$ (dashed line) as a curve tangent to the fluid velocity field vectors (black arrows) in each of the streamline’s points (two chosen are shown here as black dots, $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$). An infinitesimal increment along a streamline, $d\boldsymbol{s}$ (small blue arrow), is parallel to the local velocity vector, here normalized for convenience. Any of the streamline’s points identifies it uniquely ($s(\boldsymbol{x})\equiv s(\boldsymbol{x^{\prime}})$).
@@
with $\lambda(\boldsymbol{x})$ denoting the (curvilinear) length of the streamline, $ds$ the length of the infinitesimal increment along the streamline, and $L_{\text{ref}}$ denoting the reference length. Unless otherwise stated, we take the reference length to be equal to the length of the porous sample’s side, i.e., $L_{\text{ref}}=L$. The integration in Eq. (5) is performed over the intersection of the fluid volume $\Omega$ and plane $A$ perpendicular to the mean flow direction (see Fig. 6, top-left plot). Also, $A$ must be intersected by all the streamlines in the system. As it is ambiguous to tell the length of a looping streamline inside a recirculation zone, the integration in Eq. (5) is usually limited only to the intersection of the plane perpendicular to the mean flow direction and the percolating volume of the fluid (see Fig. 6, top-right plot)
@@
where the integration goes over the whole fluid volume $\Omega$ (see Fig. 6, bottom-left plot). We will refer to this definition as the volume-integrated tortuosity. The notation $I_{\alpha}[\>\cdot\>]$ is a shorthand for the integral of some quantity over the fluid volume or its subset, i.e.
@@
Splitting the integrals in the numerator and denominator of the volume-integrated tortuosity $T_{\Omega}$ into the contributions coming from the percolating volume $\Omega_{p}$ and the recirculation volume $\Omega_{v}$ gives
@@
According to the definition of the vortex provided earlier in this Section, the mean velocity of the vortex in steady-state flow is zero, i.e., $I[\boldsymbol{V}]_{v}=\boldsymbol{0}$, so the term $I[V_{0}]_{v}$ in the denominator can be omitted. Thus, we can rewrite Eq. (11) as
@@
Such a form of the volume-integrated tortuosity allows one to clearly distinguish between the contributions from the percolating and the recirculation zones and directly points to the sources of inequality in Eq. (10). The denominator, $I[V_{0}]_{p}$, is proportional to the Reynolds number and does not introduce any new unknown variable into our considerations.
Figure 6: Schematic representation of the three definitions of tortuosities from Eqs. (5), (7), and (8). $T_{s}^{*}$ (top left) and $T_{s}$ (top right) integrate over the surface elements $d\boldsymbol{x}$ of the intersections of an arbitrary surface AA and the corresponding fluid volumes. $T_{s}^{*}$ considers all the fluid volume $\Omega$, while $T_{s}$ – only its percolating part, $\Omega_{p}$. $T_{\Omega}$ (bottom left) integrates over the whole fluid volume $\Omega$. $T_{s}$ and $T_{\Omega}$ are used further in the present study.
II.3 Results analysis
II.3.1 Generation of streamlines and determination of the recirculation and percolating volumes
To generate the streamlines in SC samples, we numerically integrate the advection equation for the massless tracers transported in the fluid velocity field. We use the second- or third-order explicit method for the timestepping and meshless RBF-FD approximation [40] for the interpolation in space. We use the C++ implementation of meshless approximation routines provided by Medusa library [41]. We use streamlines to calculate the tortuosity and to identify the percolating/recirculation volumes in SC samples. Due to this, the initial positions of the tracers, as well as the timestep length depend on the task at hand and the considered system. The details of the streamline generation procedure and the algorithm used for the determination of percolating/recirculation volumes are delivered in Appendices C.C.4 and C.C.5, respectively. The convergence of the value of the streamline-based tortuosity $T_{s}$ with the decreasing separation of the starting points of the streamlines $h_{s}$ and advection timestep length $\Delta t_{s}$ is provided in Appendix B.
II.3.2 Calculation of the constituents of the volume-integrated tortuosity
III Results
III.1 Difference in tortuosity–Reynolds number relation for porous samples of different porosities.
We first investigate the transition from the Darcy to inertial regime for two stochastic porous samples of porosities $\phi=0.7$ and $\phi=0.9$. Top subplot of Fig. 7 shows the dependence of the tortuosity $T_{\Omega}$ defined in Eq. (8) on the Reynolds number
$Re_{k}=\frac{{\color[rgb]{0,0,0}\sqrt{k_{0}}}\langle V_{0}\rangle\phi}{\nu},\quad\langle V_{0}\rangle=\frac{I[V_{0}]}{I[1]}$ (14)
where $k_{0}$ is the intrinsic permeability . The same data are shown rescaled in the bottom subplot of Fig. 7, where by $T_{\Omega}(Re_{k}\rightarrow 0)$ we mean the value of $T_{\Omega}$ at the vanishing Reynolds number (denoted by the horizontal dashed line in the top subplot). We note that for $Re_{k}$ values beyond the investigated range, the solution went unsteady. We observe that in the Stokes regime, the tortuosity for each sample has a constant value, $1.07$ for $\phi=0.9$ and $1.175$ for $\phi=0.7$. As $Re_{k}$ increases, the tortuosity first decreases and then increases for both samples. However, the magnitude of the initial decrease and subsequent increase is considerably different for the two porosities. In the case of the higher-porosity sample, the difference between the minimum of $T_{\Omega}$ and the $T_{\Omega}(Re_{k}\rightarrow 0)$ is about $0.02$ ($0.3$ relative to the Darcy-regime value above unity, see the bottom subplot), and the increase is about $0.01$. For the lower-porosity sample, these values are $0.003$ ($0.01$ relative to the Darcy-regime value above unity) and $0.125$ ($0.67$ relative to the Darcy-regime value above unity), respectively. Such differences between the minimal and maximal values achieved by $T_{\Omega}$ in the steady-state regime were previously observed in stochastic high-porosity media [14] and in low-porosity geological samples [16]. The onset of inertial effects is also visible in the visualizations of the velocity fields, as the weaker channeling at the outlets of the samples (Fig. 8) and the emergence of the recirculation zones in the wakes of the obstacles (Fig. 9).
Figure 7: Top: the values of $T_{\Omega}$ for stochastic porous media of porosity $\phi=0.7$ (circles) and $\phi=0.9$ (squares) during the Darcy-inertial transition. The horizontal dashed lines denote the values of $T_{\Omega}$ in the limit of vanishingly small $Re_{k}$ ($Re_{k}\rightarrow 0$). Bottom: the same data as in the left subplot but rescaled by $T_{\Omega}$ at vanishingly small $Re_{k}$.
III.2 Tortuosity as the inertia onset indicator in simple cubic systems